Jump to main content or area navigation.

Contact Us

Waste Site Cleanup & Reuse in New England

Atlas Tack Corporation Superfund Site
Weekly Remedial Activities Update
Report # 17: October 30 - November 3, 2006

« Back to Atlas Tack Superfund Site Fact Sheet

1. ACTIVITIES PERFORMED DURING THE WEEK OF OCTOBER 30 - NOVEMBER 3

  • Transport and disposal (T&D) of COMM-97 soils
  • Repaired a section of earthen berm that that was damaged from the severe weather event that occurred on 28 October 2006
  • Continued to excavate soils in the NW-Fill Area
  • Began to screen soils originating from the NW-Fill Area Finalized T&D package for the disposal of oversize debris at the Bourne Landfill (1,500-cy)
  • Received analytical from the re-excavation locations along the western boundary of the Commercial and Industrial Debris Area (CID)
  • Installed tidal gauge along the northeastern boundary of the East Fill Area
  • Disassembled existing interior 8-ft high security fence in preparation of future excavation activities along the northwest corner of the NW-Fill Area. Installed an additional 70-ft section of security fence to maintain site security along the bike path.

Ongoing Site Activities:

  • Collected waste characterization soil samples from soils originating from East Fill for off site landfill profiling and disposal
  • Collected confirmatory soil samples from the floors of the NW-Fill Area
  • Continue to water the haul roads for dust control
  • Continue water handling activities
  • Continued stockpile management activities
  • Continue with air monitoring program

2. PROJECTED ACTIVITIES FOR THE WEEK OF NOVEMBER 6 - NOVEMBER 10

  • T&D Oversize Debris (1,500-cy) to the Bourne Landfill
  • Begin T&D of Subtitle-D soils to TREE
  • Return to CID and excavate area where clean-up goals have not been met: G1, F2, E2, D2, C2, A4 (see attached sketch)
  • Continue with excavation activities in the Northwest Fill Area
  • Collect characterization soil samples from soils originating from Northwest Fill Area
  • Collect confirmatory soil samples from the NW-Fill Area
  • Screen soils originating from the Northwest Fill Area

Ongoing Site Activities:

  • Continue to water the haul roads for dust control
  • Continue to collect pre and post excavation survey data of the excavations
  • Continue with stockpile management activities
  • Continue water handling activities
  • Continue with air monitoring program

3. AIR MONITORING INFORMATION

WESTON monitored for particulate (dust) during intrusive activities performed during this reporting period. WESTON recognizes EPA's level of concern for "PM10" to caution sensitive groups, including children, the elderly, and people with heart or lung disease, about exposure to coarse dust particles. The PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure.

Perimeter Air Monitoring is performed at up to (6) locations throughout a given work day. Readings are collected at 60 second intervals throughout the day and reported as a daily average.

If exceedances are identified, an increased effort of engineering controls (water suppression) will be utilized to adhere to the PM10 level of concern. If engineering controls appear to be inadequate, any dust generating activities will stop. Current dust control practices will be re-evaluated until concentrations are reduced to below the PM10 standard.

Results for 27 October – 1 November 2006 can be reviewed in the table below.

Weekly Dust Monitoring Summary
Week #17: Results for Thursday through Wednesday**

The EPA's PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure

Date Pleasant Street (µg/m3) Adams Street (µg/m3) Tripp Street (µg/m3) Egypt Road (µg/m3) Church Street (µg/m3)
10/26/2006 2 0 0 5 NSC
10/27/2006 0 0 0 7 0
10/30/2006 0 0 0 0 0
10/31/2006 3 0 14 0 0
11/1/2006 Precip Precip Precip Precip Precip

Notes:
All readings reflect the daily average calculated over the course of the workday.
All readings collected with a Thermo Electron PDR 1000.
NSC: No sampling conducted; No intrusive work conducted in that area.
Precip: No sampling conducted due to precipitation.
*: Monitoring conducted over 24-hour period.
**: No work and air monitoring occurs on weekend days

4. QUANTITY SUMMARY

Report # 16 - SUMMARY
TO DATE
ACTIVITIES TOTAL QUANTITY TOTAL QUANTITY
C.I.D. Excavation None 6,784 CY
East Fill Excavation None 9,046 CY
NW-Fill Area 3,350 CY 3,350 CY
COMM-97 Soils T&D to Crapo Landfill None 1,237 Tons
COMM-97 Soils T&D to Marion L.F. 2,220 Tons 4,160 Tons

Notes:
CY = volume of soil/debris in cubic yards

Jump to main content.